OpenStudy (anonymous):

Write 3x - 4y = 24 in slope-intercept form. a. y =3/4x + 6 b. y =3/4x - 6 c. y = 4/3x + 8 d. y = 4/3x - 8

OpenStudy (bahrom7893):

3x-4y=24 -4y=24-3x y = 3/4x - 6